OpenStudy (anonymous):

Anyone good at inverses? Find g′(−1/5), where g(x) is the inverse of f(x)=x^9/x^2+4.

OpenStudy (anonymous):

f(x)=x^9/x^2+4. (is the same as) y=x^9/x^2+4 the inverse of that is x=y^9/y^2+4 g(-1/5) means insert -1/5 in place of x in g(x) the inverse function. I hope that made sense.

OpenStudy (jdoe0001):

as shown by ehuman , find the inverse of f(x) first, then plug in it -1/5

OpenStudy (anonymous):

so is get -1/5=y^9/y^2+4 then what I do the derivative?

OpenStudy (mertsj):

Just a minute. If g is the inverse of f then f is the inverse of g. So f=g^(-1)

OpenStudy (phi):

I think they want you to use this fact \[ (f^{-1})'(y_0) = \frac{1}{f'(x_0)} \]

OpenStudy (phi):

you need to find x0 \[ \frac{x^9}{x^2+4}= \frac{1}{-5} \\ 5x^9 +x^2+4=0 \] by inspection, we (might) see x=-1 works so now the problem is to find 1/f'(-1)

OpenStudy (anonymous):

so should I plug in f(-1)?

OpenStudy (phi):

you need to find the derivative of f(x) that gives you f'(x) then evaluate at x=-1 then "flip" because you want 1/f'(-1)

OpenStudy (anonymous):

oh ok hold on one second let me see if i get right answer

OpenStudy (anonymous):

I think I did something wrong I got 0

OpenStudy (phi):

it's not zero. what do you get for the derivative?

OpenStudy (anonymous):

45x^8+2x?

OpenStudy (phi):

is \[ f(x) = \frac{x^9}{x^2+4} \] ?

OpenStudy (anonymous):

9x^8/2x

OpenStudy (phi):

you can use \[ d \left(\frac{u}{v}\right) = \frac{v \ du - u \ dv}{v^2}\]

OpenStudy (anonymous):

now I am confused were did u get that equation and what do I plug in for v and du?

OpenStudy (phi):

it is the "quotient" rule... when you have to take the derivative of a fraction where x is in the top and bottom, you use it. it says: bottom times the derivative of the top minus top times the derivative of the bottom all divided by the bottom squared.

OpenStudy (anonymous):

oh gotcha I get it now, give me a minute.

OpenStudy (anonymous):

(x^2+4)(9x^8)-(x^9)(2x) and then do i plug in the -1?

OpenStudy (phi):

yes, but remember all of that is divided by (x^2+4)^2

OpenStudy (anonymous):

Oh yea lol, alright let me plug and see if I get answer.

OpenStudy (anonymous):

I got 44.92

OpenStudy (anonymous):

did you get 25 on the bottom?

OpenStudy (phi):

the bottom is 25 the top is 43 so f'(-1)= 43/25 and 1/f'(-1) = 25/43 which is the answer

OpenStudy (anonymous):

Hmm I hate 1.72 and typed it in and got it wrong and then I saw u flipped it but it still did not work.

OpenStudy (phi):

I would enter is 25/43

OpenStudy (anonymous):

Oh that worked, maybe it only takes fractions which makes sense for calculus! Thanks for all the help!